Assessment plots Due in 14 hours, 6 m Here is a set of sample data 1 15 26 38 62 67 79 84 93 96 100 Identify the 5 number summary (min, Q1, median, Q3, max) Hint: HelpG Video Textbook Points possible: 1 This is attempt 1 of 3. 3S Submit ctionaskip&ito 3
Answer #1

Here n = 11

So median is the (n+1)/2 th value in the arrange data set.

so median = (11 + 1)/2 th value = 6th value = 67

The first quartile is the median of the first five values = (5+1)/2 th value = 3rd value = 26

The 3rd quartile is the median of the last five values = { 6 + (5+1)/2 }th value = 9th value = 93

Minimum = 1

maximum = 100

Look the following image:

1 This is the minimum value 15 26 this is the first quartile value 38 62 67 This is the median value 79 84 93 This is the 3rd quartile value. 96 3 4 5 10 12 100 This is the maximum value 12

So correct choices are as follows:

Min Q1 median Q3 max
1 26 67 93 100